OpenStudy (anonymous):

If t = -2π/3, evaluate (if possible) sin t, cos t, tan t

OpenStudy (anonymous):

just type it into your calculator. sin(-2pi/3) etc. and it should give you and answer

OpenStudy (anonymous):

i know but work needs to be shown.

OpenStudy (anonymous):

using an unit circle. . i need to find what point the terminal side lands on.

OpenStudy (anonymous):

sin (-theta) = - sin theta cos (-theta) = cos theta tan (-theta) = -tan (theta) Use this and simply find the values of sin (2π/3) , cos (2π/3) and tan (2π/3) and attach the appropriate sign to it.

OpenStudy (anonymous):

so the points can either be. . (0,1) (0,-1) (-1,0) (1,0)

OpenStudy (anonymous):

is there a way ou can change -2π/3 into degrees

OpenStudy (anonymous):

See, plot a unit circle. -2π/3 will land in 3rd quadrant. So simply make an angle of π/3 with the negative x-axis and draw perpendiculars to the -ve Y and -ve X axis. Each perpendicular will give you the value of Sin -2π/3 (= -root3/2) and cos -2π/3(= -1/2) respectively. If you are wondering why the cos value has a negative sign, it is because although cos(-theta) = cos theta, but since the angle here is 2π/3 (and cos is negative in the second quadrant) we add another negative sign.

OpenStudy (anonymous):

to change -2π/3 to degrees simply multiply it by 180/π. You'll get the value in degrees (=-120)

OpenStudy (anonymous):

thanks. so how do i find csc t , sec t , & cot t

OpenStudy (anonymous):

Now that you have the values of sin t cos t and tan t. Just use their reciprocals to find the values of cosec t sec t and cot t